OpenStudy (anonymous):

A cone has a volume of 20 cubic inches and the area of the base is 15 square inches,What is the height of the pyramid?

OpenStudy (anonymous):

lol where the pyramid came from?? in case of height of cone 1/3*area of base*height = volume 1/3*15*h=20 (let h be the height of cone) you can now solve to have value of h.

OpenStudy (amistre64):

isnt a cone 1/3 the volume of it cuylindar counterpart?

OpenStudy (anonymous):

@theyatin sorry wrong given here is it : A pyramid has a volume of 20 cubic inches and the area of the base is 15 square inches,What is the height of the pyramid?

OpenStudy (anonymous):

even then same formula is applied in both cases. . . same formula is used in pyramid too pyramid is cone with square base lol :P

OpenStudy (anonymous):

@Faye281

OpenStudy (anonymous):

ok thanks